OpenStudy (anonymous):

Fan and medal 2x^2+x=3 simplify

OpenStudy (solomonzelman):

you are solving over here, not simplifying, since you have ` = ` in your equation.

OpenStudy (solomonzelman):

\(\large\color{red}{ \rm 2x^2+x=3 }\) \(\large\color{red}{ \rm 2x^2+x-3=0 }\) \(\large\color{red}{ \rm (x-1)(2x+3)=0 }\) \(\large\color{red}{ \rm x= ? }\)

OpenStudy (anonymous):

Is it (1,-3/2)

OpenStudy (solomonzelman):

yes 1 and -3/2